OpenStudy (anonymous):

Help with finding f^-1 (x)

OpenStudy (anonymous):

f(x)= \[\sqrt[3]{x-7}\]

OpenStudy (johnweldon1993):

you would need to switch the x's and y's in your equation

OpenStudy (anonymous):

step1; cube both sides

OpenStudy (johnweldon1993):

and then solve for y

OpenStudy (anonymous):

step2; isolate x

OpenStudy (anonymous):

step3; replace y with x and vice versa

OpenStudy (johnweldon1993):

so f^-1 would look like y = cubed root of x - 7 x = cubed root of y -7 so you would cube both sides x^3 = y-7 add 7 to both sides y = x^3 + 7

OpenStudy (anonymous):

so f^-1 (x) would be x^3+7?

OpenStudy (anonymous):

Once you replace y with x and x with y and solve for y, you have found the inverse.

OpenStudy (johnweldon1993):

that would be correct.....as @calmat01 has said...one you have swtiched the x's and y's...then solved for y again..you have found the inverse